OpenStudy (anonymous):

An equation is shown below: 5(2x - 8) + 15 = -15 Write the steps you will use to solve the equation and explain each step.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

not really :S

OpenStudy (anonymous):

first we have to do whats in the parenthesis

OpenStudy (anonymous):

right?

OpenStudy (skullpatrol):

How about subtracting 15 from both sides first, then divide both sides by 5?

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

so it would be -30 right? on one side

OpenStudy (skullpatrol):

Yup.

OpenStudy (anonymous):

divided by 5 = -6

OpenStudy (skullpatrol):

Yup

OpenStudy (anonymous):

so now its (2x - 8) = -6

OpenStudy (anonymous):

?

OpenStudy (skullpatrol):

Correct

OpenStudy (anonymous):

so what do we do now

OpenStudy (skullpatrol):

Remove the parentheses

OpenStudy (anonymous):

ok so 2x - 8 = -6

OpenStudy (skullpatrol):

Correct

OpenStudy (anonymous):

next... lol

OpenStudy (skullpatrol):

What is your goal?

OpenStudy (anonymous):

solve the equation

OpenStudy (skullpatrol):

Isolate x means get x alone on one side of the equation, right?

OpenStudy (anonymous):

i guess yeah

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

Any ideas? Recall what we did earlier :-)

OpenStudy (anonymous):

add 8?

OpenStudy (anonymous):

it would = 2

OpenStudy (skullpatrol):

Yes, to both sides

OpenStudy (anonymous):

so then it would be 2x = 2

OpenStudy (skullpatrol):

Yup

OpenStudy (anonymous):

so then?

OpenStudy (skullpatrol):

What value of x will make 2x = 2 a true statement?

OpenStudy (anonymous):

1

OpenStudy (skullpatrol):

Correct, now go back to the original equation and check it.

OpenStudy (anonymous):

so i plug in 1 for x

OpenStudy (skullpatrol):

Yup

OpenStudy (anonymous):

5(2*1 - 8) + 15 = -15

OpenStudy (anonymous):

how do i check it? idk where to start

OpenStudy (skullpatrol):

Inside the parentheses

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

2 - 8?

OpenStudy (anonymous):

-6

OpenStudy (skullpatrol):

Yup

OpenStudy (anonymous):

so then is that it or what lol

OpenStudy (skullpatrol):

5(-6) + 15 = ?

OpenStudy (anonymous):

-30 + 15 = -15

OpenStudy (skullpatrol):

Is -15 = -15 a true statement?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

done? thanks!

OpenStudy (skullpatrol):

You have just proven that x = 1 is the solution :D

OpenStudy (anonymous):

lol thanks

OpenStudy (skullpatrol):

Thanks for trying to learn.